# Sentiment and Social Signals in the Climate Crisis: A Survey on Analyzing Social Media Responses to Extreme Weather Events

### v1

@@ 1. Introduction
Figure 1. Overview of the climate sentiment pipeline linking extreme events to social media reactions, analyzed through data, models, and their societal impact. Our survey follows this structure, reviewing existing literature across each stage of the pipeline.
[…]
Diagram showing the flow from climate events to social media reactions, followed by stages such as data collection, modeling, and real-world applications like policy and health monitoring.
@@
This survey provides a thorough overview of sentiment analysis techniques used in social media related to weather and climate change events  [^1] To support transparency and reproducibility, an anonymous GitHub repository is maintained at https://github.com/whisperoghost/ClimateSentiment, where the the list of referenced papers is available and open for community contributions.. While there are general reviews available, very few specifically address the unique challenges, methodologies, and insights that emerge at the intersection of climate disasters and public sentiment (Vivion et al., 2024; Ahmad, 2022; Sayigh, 2023). To frame our discussion, we will use the 2025 Los Angeles forest fires as a case study. This large-scale wildfire sparked intense online discourse, characterized by grief, outrage, political polarization, and systemic critique (Farahmand et al., 2025; Gimello, 2025).
@@
We examine methods for real-time data collection and annotation, including hashtag tracking, social media scraping, and weak supervision, which are used to create labeled datasets during rapidly evolving climate events (Olteanu et al., 2015; Al-Saqaf and Berglez, 2019; Fernandez et al., 2016). Special attention is given to the challenges of annotating sarcasm, emotional nuance, and cross-cultural expressions of emotion in high-stakes and politically sensitive contexts (Mäkelä, 2024; Pranto, 2024; Pupneja et al., 2023). We also examine current evaluation strategies and datasets, highlighting key research gaps: the absence of multilingual and multimodal corpora, insufficient real-time modeling of sentiment evolution, and the inadequate representation of misinformation-aware sentiment systems (Kokoschka et al., 2024; Yosipof et al., 2024; Vivion et al., 2024; Lakzaei et al., 2024). A primary motivation for this survey is the increasing prevalence of AI-generated misinformation during climate emergencies. (Yosipof et al., 2024; Gimello, 2025; Michail et al., 2024). Events like the 2025 LA forest fires revealed how synthetic images and videos, such as fake depictions of burning Hollywood landmarks, can mislead the public. (Chappell, 2025; Afshar, 2025; Gimello-Mesplomb, [n. d.]).
@@ 4.3. Deep Learning Approaches
Convolutional Neural Networks (CNNs): Originally designed for image processing, CNNs have been adapted for text classification by treating sentences as matrices of word embeddings. They are particularly good at capturing local n-gram patterns (Kim et al., 2022; Kareem et al., 2021; Farhangmehr et al., 2025).
@@ 4.4. Transformer-Based Models
ClimateBERT: A domain-adapted BERT model fine-tuned on climate change-related texts, improving relevance for environmental discourse (Webersinke et al., 2021; Garrido-Merchán et al., 2023; Schimanski et al., 2023).
@@ 4.5. LLMs for Sentiment Analysis
The methodological landscape for sentiment analysis in climate discourse has evolved significantly over the last decade. The choice of method depends on the nature of the task, the quality and size of available data, the urgency of analysis, and the interpretability or ethical constraints of the application. In the following sections, we delve into the practical deployment of these models, supported by case studies and performance benchmarks.
@@ Acknowledgements.
The author would like to thank the faculty and colleagues at the School of Computing and Augmented Intelligence, Arizona State University, for their valuable feedback and support during the development of this work. Special thanks to the SHaDE Lab for ongoing inspiration in interdisciplinary climate research, and to the wider NLP and computational social science communities for their contributions to open tools, datasets, and DMML lab for discussions that shaped this survey. Automated tools, including GPT-based models, were used to assist with grammar refinement and minor language editing for clarity and fluency.

### v4

@@ 1. Introduction
Figure 1. Overview of the climate sentiment pipeline linking extreme events to social media reactions, analyzed through data, models, and their societal impact. Our survey follows this structure, reviewing existing literature across each stage of the pipeline. :
@@
This survey provides a thorough overview of sentiment analysis techniques used in social media related to weather and climate change events  [^1] The full and extended literature review—including all references analyzed and categorized beyond those cited in this manuscript—is available at the project’s GitHub repository: https://github.com/pouyashaeri/ClimateSentiment. While there are general reviews available, very few specifically address the unique challenges, methodologies, and insights that emerge at the intersection of climate disasters and public sentiment (Vivion et al., 2024; Ahmad, 2022; Sayigh, 2023). To frame our discussion, we will use the 2025 Los Angeles forest fires as a case study. This large-scale wildfire sparked intense online discourse, characterized by grief, outrage, political polarization, and systemic critique (Farahmand et al., 2025; Gimello, 2025).
@@
We examine methods for real-time data collection and annotation, including hashtag tracking, social media scraping, and weak supervision, which are used to create labeled datasets during rapidly evolving climate events (Olteanu et al., 2015; Al-Saqaf and Berglez, 2019; Fernandez et al., 2016). Special attention is given to the challenges of annotating sarcasm, emotional nuance, and cross-cultural expressions of emotion in high-stakes and politically sensitive contexts (Mäkelä, 2024; Pranto, 2024; Pupneja et al., 2023). We also examine current evaluation strategies and datasets, highlighting key research gaps: the absence of multilingual and multimodal corpora, insufficient real-time modeling of sentiment evolution, and the inadequate representation of misinformation-aware sentiment systems (Kokoschka et al., 2024; Yosipof et al., 2024; Vivion et al., 2024; Lakzaei et al., 2024). A primary motivation for this survey is the increasing prevalence of AI-generated misinformation during climate emergencies. (Yosipof et al., 2024; Gimello, 2025; Michail et al., 2024). Events like the 2025 LA forest fires revealed how synthetic images and videos, such as fake depictions of burning Hollywood landmarks, can mislead the public (Chappell, 2025; Afshar, 2025; Gimello-Mesplomb, [n. d.]).
@@ 4.3. Deep Learning Approaches
Convolutional Neural Networks (CNNs): Originally designed for image processing, CNNs have been adapted for text classification by treating sentences as matrices of word embeddings. They are particularly good at capturing local n-gram patterns (Kim et al., 2022b; Kareem et al., 2021; Farhangmehr et al., 2025).
@@ 4.4. Transformer-Based Models
ClimateBERT: A domain-adapted BERT model fine-tuned on climate change-related texts, improving relevance for environmental discourse (Webersinke et al., 2021; Garrido-Merchán et al., 2023a; Schimanski et al., 2023).
@@ 4.5. LLMs for Sentiment Analysis
The methodological landscape for sentiment analysis in climate discourse has evolved significantly over the last decade. The choice of method depends on the nature of the task, the quality and size of available data, the urgency of analysis, and the interpretability or ethical constraints of the application. In the following sections, we delve into the practical deployment of these models, supported by case studies and performance benchmarks. A comprehensive comparison of sentiment analysis approaches applied to climate-related social media is presented in Table 1, organizing key literature by methodology, performance, key findings, and limitations.
Table 1. Comprehensive comparison of sentiment analysis methodologies for climate-related social media
Study Method Data/Event Performance Key Findings and Limitations
Lexicon-Based Methods
Rosenberg et al. (Rosenberg et al., 2023) VADER, TextBlob Twitter/Climate SDGs VADER: 56%, TextBlob: 46% accuracy VADER better; keyword bias remains
Mohamad Sham et al. (Sham and Mohamed, 2022) Hybrid (TextBlob + LR) Twitter/Climate discussions 75.3% F1-score Hybrid outperforms individual models
Mendon et al. (Mendon et al., 2021) TextBlob Twitter/Kerala Floods 9% neg, 71% neutral, 20% pos Combined with clustering; polarity limitations
Amangeldi et al. (Amangeldi et al., 2024) PMI, NRCLex Twitter, Reddit, YouTube 65% accuracy PMI outperforms VADER; domain tuning needed
Traditional Machine Learning Methods
Kim et al. (Kim et al., 2022a) SVM, NB, LR, RF, XGBoost News/ Environmental categories 73% F1-Score ML models with term weighting; lacks sentiment context
Sabri et al. (Sabri et al., 2025) SVM Twitter/Climate awareness 91% accuracy, 90% F1-score SVM with TF-IDF; focused on awareness detection
Reyes-Menendez et al. (Reyes-Menendez et al., 2018) SVM (MonkeyLearn) Twitter /#WorldEnvironmentDay 72.4% accuracy SVM maps sentiment to SDGs; sarcasm misclassification
Thenmozhi et al. (Thenmozhi et al., 2024) SVM with TF-IDF Twitter/Climate debates 76.3% accuracy, 70.5% F1-score SVM captures nuanced sentiment
Deep Learning Approaches
Rumlawang et al. (Rumlawang et al., 2025) LSTM Twitter/Climate sentiment (Indonesia) 85.6% accuracy, 84% F1-score LSTM captures temporal patterns; small dataset
Wu et al. (Wu et al., 2023) BiLSTM Weibo/Climate topics $\sim 90$% accuracy BiLSTM reveals spatio-temporal trends; China-specific
Rumlawang et al. (Rumlawang et al., 2025) LSTM Twitter/Climate change sentiment (Indonesia) 60% accuracy, 59% F1-score LSTM outperforms ML baselines; struggles with neutral vs negative distinction
El Barachi et al. (El Barachi et al., 2021) BiLSTM Twitter/Real-time sentiment 89.3% accuracy Real-time tracking; class imbalance issues
Marjani et al. (Marjani et al., 2024) CNN-BiLSTM Social + Satellite/Wildfires 73% F1-score train, 58% F1-score val Social-physical signal fusion; moderate temporal accuracy
Transformer-Based Models
Leippold et al. (Leippold et al., 2022) Climate BERT News/Climate risk texts 35.7% error reduction vs BERT Domain pretraining boosts performance
Ong et al. (Ong et al., 2024) RoBERTa, DistilBERT Twitter/ Sustainable investing 88% accuracy Handles sentiment and topics; financial bias risk
Garrido-Merchan et al. (Garrido-Merchán et al., 2023b) BERT Financial Reports 85% F1-score Fine-tuned for climate risk; finance-focused
Krishnan and Anoop (Krishnan and Anoop, 2023) Climate BERT Twitter/Global climate 90.2% accuracy, 85.5% F1-score Social sentiment focused; small dataset
Large Language Models
Anderson et al. (Anderson et al., 2024) GPT-2 Twitter/ Sustainability topics 84.3% accuracy GPT-2 beats ML baselines; limited fine-tuning
Jeng et al. (Jeng et al., 2025) GPT-4o News/Climate journalism Behavioral improvement on intent GPT-4o for emotional framing; lacks standard metrics
Fan and Xu (Fan and Xu, 2025) GPT-4o, Claude 3.5, Gemma, Llama 3/3.3 Twitter, Reddit/Climate change 52-68% neutral responses; significantly lower emotional intensity (p<<0.001) LLMs demonstrate emotional moderation in climate discussions
Linardos et al. (Linardos and Kermanidis, 2025) GPT-4o-mini Twitter/Disaster response 70-86% accuracy by category; 66-87% F1-score LLM aids tweet understanding; used with ML clustering
The evolution from lexicon-based methods (65-75% accuracy) to Transformer-based methods and LLMs (85-89% accuracy) shows clear performance progression. Lexicon methods offer interpretability but struggle with context. Traditional ML requires feature engineering but provides efficiency. Deep learning captures complex patterns but needs large datasets. Transformers achieve strong contextual understanding, while LLMs offer flexibility through prompting but raise reliability concerns.
@@ 5.2. Comparison with Past Events
To highlight temporal shifts in public climate discourse, we compare sentiment patterns across major climate events. Table 2 summarizes dominant emotions and key characteristics for each event, reflecting changes in emotional tone and the increasing influence of AI-generated content.
Table 2. Comparative analysis of sentiment patterns across major climate events
Event Year Dominant Emotions Key Characteristics
Australian Bushfires 2019-20 Sadness, grief, solidarity Global empathy, wildlife focus, minimal AI
Hurricane Ida 2021 Fear, frustration, resilience Infrastructure failure, regional discourse
European Heatwaves 2022-23 Anger, irony, anxiety Multilingual content, satirical responses
LA Forest Fires 2025 Fear, anger, helplessness Extensive synthetic media, political blame
@@ Acknowledgements.
The author thanks the faculty and colleagues at the School of Computing and Augmented Intelligence, Arizona State University, for their feedback and support. Special appreciation goes to the SHaDE Lab for inspiring interdisciplinary climate research, the DMML Lab for insightful discussions, and the NLP and computational social science communities for open tools and datasets. Automated tools, including GPT-based models, were used for minor language and grammar refinements.